OpenStudy (genericnoodles):

Algebra 1-Factoring Having trouble with factoring, the equation is below. 15v^2+99v+120

OpenStudy (anonymous):

you can take out a common factor of 3 first. That'll help a little bit.

OpenStudy (anonymous):

do you have a graphing calculator? that's usually what I use to help me with tough ones like that.

OpenStudy (genericnoodles):

I do not have one, is dividing by 3 the last step?

OpenStudy (anonymous):

not quite. you still have to factor the binomial, but now you have 5v^2 +33v+40

OpenStudy (genericnoodles):

Shouldn't the final answer be 3(5v2+33v+40)?

OpenStudy (anonymous):

Factor out 3 and you are left with: 3(5v^2+33v+40) The only numbers that multiply together to equal 5v^2 and both have v in them are: 5v and v Now set up your parenthesis. Everything is positive, so you know that both are plus: (5v+?)(v+?) The question marks are unknown numbers. They must multiply together to equal 40, but add together to equal 33 (one must be multiplied by 5). What 2 numbers multiplied equals 40? 20 and 2, 10 and 4, 8 and 5. One of these numbers * 5 + the other number must equal 33. 20+2(5)=30, that's not it! 10+4(5)=30, that's not it either! 8+5(5)=33, ding ding ding! So the 2 numbers are 8 and 5, but the 5 is in the parenthesis that does not include 5v because you have to multiply the two. So you're left with: \[3(5v+8)(v+5)\]

OpenStudy (genericnoodles):

Thank you a lot!